OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4?

OpenStudy (tkhunny):

You will need the Point-Slope form. Can you quote it, please?

OpenStudy (anonymous):

y-y1=m(x-x1)?

OpenStudy (tkhunny):

Very good. Now, the point we need is just given to us. The only difficulty is the slope. The slope information is in the perpendicular line. If the slope of our line is perpendicular to y = -2x + 4, what is the slope of our line? There is a rule about that. This problem is designed to see if you now the rule.

OpenStudy (anonymous):

...But I don't know the rule. :(

OpenStudy (tkhunny):

Well, then here it is. If two lines are perpendicular (and neither is horizontal), their slopes multiply to -1. If one line has a slope of -2, we have -2 * m = -1 and lines perpendicular to it have a slope of 1/2. (-2)*(1/2) = -1 If one line has a slope of 7/3, we have (7/3) * m = -1 and lines perpendicular to it have a slope of -3/7. (7/3)*(-3/7) = -1 Can you now solve the given problem?

OpenStudy (anonymous):

Yes, I solved it and got it right. Thanky you.